OpenStudy (anonymous):

5x+3y = -9 3x-5y=-9.....how do i solve this using elimination?

OpenStudy (anonymous):

5x+3y = -9 => x = ( -9-3y/5 Putting this value of x in equation 3x -5y=-9 you will get the value of y .

OpenStudy (anonymous):

im still confused,can you explain a little more

OpenStudy (anonymous):

Multiply 1st equation by 3 and the other by 5, we get, \[15x + 9y = -27\] \[15x - 25y = -45\] Then Subtract one and second equation: \[34y = 18\] And y = \[18/34 = 9/17\] And therefore x = \[-36/17\].. Now understood??

OpenStudy (anonymous):

sorry that was substitution and not elimination method

OpenStudy (anonymous):

i fully understand it now! thanks

OpenStudy (anonymous):

No thanks, just knowledge... And you know what, If you have knowledge, Let others light their candle on it...